OpenStudy (anonymous):

If y = (x + 3)2, and the value of x=5, what is the value

OpenStudy (anonymous):

y=(5+3)2 y=(8)2 y=16

OpenStudy (anonymous):

now to check your answer, 16=(5+3)2 16=(8)2 16=16